OpenStudy (anonymous):

the # of pages mara can read varies directly with time. she can read 15 pages in 20 minutes . how many pages can she read????

OpenStudy (anonymous):

wat does int mean???

OpenStudy (anonymous):

n = pages t= time Directly proportional uses the equation n=kt or k=n/t First solve for k n=kt 15=k*20 k=15/20 the number of pages she reads is n=(15/20)*t and int is me

OpenStudy (anonymous):

but wat does it stand for?

OpenStudy (anonymous):

uh. I dont know. I just picked it. haha

OpenStudy (anonymous):

ok then im gonna post a ? if u wanna help me out.......

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

k gimmme a sec

OpenStudy (anonymous):

its posted